Rainluke01:

what number does 3 and 26 go into but can be dived by 5 evenly?

Pixel:

i think it is 130

Pixel:

26 x 5 = 130

Pixel:

but i would check it

Vocaloid:

to find the least common multiple of 3,26, and 5, write out the prime factorization for each number 3 = 3 26 = 2 * 13 5 = 5 in this case, since each prime factor only appears once we simply multiply 3 * 13 * 2 * 5 = your ans